OpenStudy (anonymous):

integrate[sin(theta)^2] using integration by parts. The final answer must be in the form (-1/a) * sin(B *theta) * cos(D * theta) + (1/F)* theta + C

OpenStudy (anonymous):

right sin(theta )^2 as [sin(theta).sin(theta)]

OpenStudy (anonymous):

then use by parts

OpenStudy (anonymous):

If I do that I get this: -\[\sin \Theta \cos \theta - \int\limits_{}^{}-\cos^2\theta d \theta\] Which isn't any easier to solve for.

OpenStudy (anonymous):

Or is it? I don't think it is.